On December 31, 2005, Merlin Company had outstanding 400,000 shares of common stock and 40,000 shares of  
8% cumulative preferred stock (par $10). On February 28, 2006, Merlin issued an additional 36,000 shares of common stock.  
On September 1, 2006, 9,000 shares were retired. Net income was $900,000. A 10% stock dividend was
declared and distributed on July 1, 2006
What is the basic EPS for the year ended December 31, 2006 for Merlin Company?
Answer #1

Weighted average shares = (400000*1.1*2/12)+(436000*1.1*4/12)+(436000*1.1*2/12) +(470600*4/12) = 470000

Preferred dividend = 40000*10*8% = 32000

EPS = (Net income-Preferred dividend)/Share outstanding = (900000-32000)/470000 = 1.85 per share